### v1

# Anisotropic resistivity of a $p$-wave magnet candidate CeNiAsO

Abstract
Certain noncollinear antiferromagnets host momentum-dependent spin-splitting bands with odd-parity $p$-wave symmetry, termed $p$-wave magnets. Metallic $p$-wave magnets offer a unique platform for spintronic applications, yet experimental realizations remain scarce. CeNiAsO, featuring a tetragonal structure and developing a commensurate coplanar noncollinear antiferromagnetic order at low temperatures, has been theoretically proposed as such a magnet. Here, we show that its resistivity exhibits strong two-fold in-plane anisotropy — a key signature of its proposed $p$-wave magnetism. Reversible and nonvolatile switching between high- and low-resistance states is achieved by alternating the in-plane magnetic field direction, indicating potential for field-controlled memory devices. Our results not only provide critical experimental evidence for $p$-wave magnetism in CeNiAsO but also open new avenues for high-performance antiferromagnetic spintronics based on resistivity anisotropy.
The concept of altermagnetism has recently been proposed to achieve momentum-dependent, nonrelativistic spin-splitting band structures in certain antiferromagnets (AFMs). Initial studies focused on collinear AFMs exhibiting even-parity spin polarized energy iso-surfaces with $d$-, $g$-, or $i$-wave symmetry [1, 2]. Materials like MnTe, CrSb, RuO2, and KV2Se2O have demonstrated characteristic altermagnetic behaviors [3, 4, 5, 6]. Subsequent theoretical studies revealed that $p$-wave symmetry with odd parity can be realized in noncollinear AFMs, termed $p$-wave magnets [7, 8, 9, 10]. Crucially, $p$-wave magnets preserve $\mathcal{T}t$ symmetry combining time-reversal $\mathcal{T}$ and a lattice translation $t$, contrasting with altermagnets where this symmetry is broken. Since the momentum-dependent collinear spin polarization is orthogonal to the coplanar spins in the real space, they have been characterized as antialtermagnets [8].
Although real-space cancellation of spin-splitting bands impedes spin current generation, $p$-wave magnets remain compelling platforms for exploring novel phenomena, especially via proximity coupling with superconductivity [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Experimentally, $p$-wave behavior has been observed in the type-II multiferroic, NiI2 [22], though its insulating nature hinders applications. The intermetallic compound Gd3(Ru1-xRhx)4Al12 ($x\approx$ 0.05) also shows $p$-wave behavior [23], yet its large anomalous Hall effect demonstrates the existence of $\mathcal{T}t$ symmetry breaking, contradicting $p$-wave magnetism’s core requirement. Moreover, both possess hexagonal crystal structures with helical spin configurations. This contrasts sharply with CeNiAsO, theoretically proposed as a $p$-wave magnet, which exhibits both a tetragonal structure and a commensurate Néel-type spin order [7].
CeNiAsO is a structural analogue of the 1111-type iron-based superconductors, crystallizing in the tetragonal ZrCuSiAs-type structure with space group $P4/nmm$ (No. 129) [24]. Its lattice constants are comparable to those of cuprates and iron-based superconductors, which may favor coupling with high-temperature superconductivity. This compound undergoes two successive AFM transitions at $T_{N1}$ and $T_{N2}$ with magnetic moments on Ce ions [24, 25, 26]. Both pressure and phosphorus doping suppress the AFM orders and induce a quantum critical point [27]. These results highlight the heavy-fermion behavior of the electronic system. Theoretical work further proposes CeNiAsO as a promising $p$-wave candidate [7], originating from its low-temperature non-collinear, coplanar magnetic structure [Fig. 1(a)]. This structure possesses a nonrelativistic spin symmetry described by $C_{2\perp}||t$], which is equivalent to the magnetic symmetry $\mathcal{T}t$ in coplanar magnets, where $C_{2\perp}$ denotes a two-fold rotation in spin space around the $c$-axis, and $t$ represents lattice translation. Notably, all prior studies used polycrystalline samples, whereas single crystals are essential for investigating its $p$-wave magnetic properties, particularly the large spontaneous anisotropy of in-plane resistivity [7].
In this work, we successfully grew single crystals of CeNiAsO using the self-flux method and observed a large two-fold in-plane resistivity anisotropy in the low-temperature noncollinear Néel state, consistent with theoretical predictions for $p$-wave magnetism [7]. This anisotropy emerges exclusively upon magnetic field-controlled domain selection. The observed resistivity anisotropy reaches up to approximately 35 %, which highlights the promising potential for spintronic applications. Reversible and nonvolatile resistivity switching is achieved by alternating the in-plane magnetic field direction. Our results not only provide compelling experimental evidence for the $p$-wave magnetic nature of CeNiAsO but also establish a foundation for domain-engineered antiferromagnetic spintronics, where resistance states can be effectively controlled through manipulation of the magnetic domain population.
CeNiAsO single crystals were grown via a two-step process. First, polycrystalline samples were synthesized by solid-state reaction as reported previously [24]. Second, single crystals were grown by the self-flux method using NaAs as flux. Polycrystalline CeNiAsO and NaAs powders were loaded into an alumina crucible at a 1:2 mass ratio. The crucible was first sealed in an Nb tube and then in an evacuated quartz tube. The mixture was heated to $1300^{\circ}\mathrm{C}$ over 10 hours, held for 10 hours to ensure complete melting, then slowly cooled to $1200^{\circ}\mathrm{C}$ at $1^{\circ}\mathrm{C/h}$ before furnace shutoff. Typical crystals have a square shape with the size of approximately 0.4 $\times$ 0.4 $\times$ 0.15 mm3. Single-crystal X-ray diffraction (SCXRD) was used to verify crystal structure and chemical composition. The crystalline axes were checked by a Laue diffractometer. Specific-heat and resistivity measurements were performed on a physical property measurement system (Quantum Design). Magnetic susceptibility was measured on a magnetic property measurement system (Quantum Design).
Figure 1: (a) Sketch of the commensurate Néel (upper panel) and incommensurate SDW (lower panel) in-plane magnetic structures of CeNiAsO for $T<T_{N2}$ and $T_{N2}<T_{N1}$, respectively [25]. Only cerium atoms are shown for simplicity. The red and blue dashed lines define the nuclear and Néel-order magnetic in-plane unit cells, respectively. While $a_{T}$ and $b_{T}$ lattice constants are the same in the tetragonal symmetry, lattice constants $a_{M}$ and $b_{M}$ in the magnetic unit cell are equal to $2a_{T}$ and $b_{T}$, respectively. This results in two domains (D1 and D2) for each AFM order. (b) Temperature dependence of in-plane resistivity. The inset shows its first derivative at low temperatures where $T_{N1}$ and $T_{N2}$ are labeled. (c) The temperature dependence of specific heat $C/T$ below 14 K at 0 and 7 T. (d) Temperature dependence of magnetic susceptibility $\chi=M/H$ for the field $H$ parallel to $a$ and $c$ axes, respectively. The blue line is the Curie-Weiss fit result for the $\chi_{a}$. The inset shows the $M$-$H$ loop at 3 K with $H//a$.
[…]
The presence of these magnetic phases is clearly revealed in our transport measurements. Figure 1(b) displays the temperature dependence of in-plane resistivity at 0 T, with its first derivative shown in the inset. $T_{N1}$ and $T_{N2}$ are 8.8 and 6.2 K, respectively. Figure 1(c) presents the low-temperature specific heat versus temperature, where a 9-T field has almost no effect on the specific heat. These results confirm two AFM transitions consistent with prior polycrystalline studies [24]. Figure 1(d) shows temperature-dependent magnetic susceptibility, revealing strong magnetic anisotropy for $H//a$ and $H//c$ from 300 K down to 2 K. The high-temperature $H//a$ data ($\chi_{a}$) can be fitted by the Curie-Weiss function, $\chi=\chi_{0}+C/(T-\theta_{W})$, where $\chi_{0}$, $C$, and $\theta_{W}$ are the temperature-independent background, Curie constant, and Weiss temperature, respectively. The fitting yields $\theta_{W}$ = -23.8 K and $\mu_{eff}$ = 2.11 $\mu_{B}$, close to those reported for polycrystalline samples. The weak temperature dependence of $\chi_{c}$ prevents reliable Curie-Weiss fitting.
Figure 2: (a) Field dependence of resistance at 1.8 K. The current is applied along the lattice $a$-axis. $x$ and $y$ directions are defined as parallel and perpendicular to the current direction, respectively. The arrows indicate the field application process. (b) Angular dependence of resistance at 1.8 K. The arrows indicate ascending or descending angle scans. (c) Field dependence of resistance at 7 K. (d) Angular dependence of resistance at 7 K. The resistance values in (b) and (d) have been shifted by a constant at each field for better presentation.
It has been theoretically proposed that anisotropic resistivity serves as a fingerprint of a $p$-wave magnet, arising from a spin-split Fermi surface with odd parity [7]. Consequently, for CeNiAsO below $T_{\mathrm{N2}}$, a significant resistivity anisotropy is expected along the $a_{M}$- and $b_{M}$-axes. However, in a zero-field-cooled (ZFC) sample, the degenerate domains D1 and D2 may coexist with equal probability, leading to an effective averaging of the direction-dependent resistivity and resulting in isotropic transport behavior. To suppress this domain averaging and unveil the intrinsic resistivity anisotropy, we apply an in-plane magnetic field—along either the $x$- or $y$-axis—to lift the degeneracy between D1 and D2 by introducing an energy splitting between the two domains.
Figure 2(a) shows the field dependence of resistance at 1.8 K with the field direction parallel to either $x$ or $y$ direction. For $H//y$, the resistance increases rapidly from the ZFC value $R_{ZFC}$(0T)] with increasing field without saturation up to 9 T. Afterwards, the sample remains in a high-resistance state, corresponding to the resistivity along the $a_{M}$-axis, indicating that domain D1 is dominant in this case [Fig. 1(a)]. Intriguingly, butterfly-shaped hysteresis emerges, and the field-trained zero-field resistance $R_{FT}$(0T)] (i.e., zero-field resistance after field sweeping up to 9 T and then down to 0 T) does not depend on field application history anymore. This behavior may be attributed to a slight remodulation of the relative concentrations of domains D1 and D2 in a sample where D1 is initially dominant. A complementary behavior is observed for $H//x$, where the resistance initially decreases with increasing field and then the sample stays in a low-resistance state. This state corresponds to the resistivity along the $b_{M}$-axis, indicating that domain D2 becomes dominant under these conditions. We note that such hysteresis is not associated with spin canting as demonstrated in Fig. 1(d), where magnetization depends linearly on magnetic field without detectable hysteresis. This behavior is consistent with our assumption that the magnetic field only remodulates the domain populations or alignments, rather than tilting the magnetic moments.
To better understand the anisotropic resistance, we measured the angular-dependent magnetoresistance (AMR) with the field rotating within the $ab$ plane. Figure 2(b) shows the results at 1.8 K. All curves exhibit 180-degree periodicity, confirming the in-plane two-fold symmetry of resistivity anisotropy. At 9 T, the AMR follows a nearly squared functional form rather than trigonometric functions, indicating that the anisotropy is not associated with the projection of the magnetic field. Significant hysteresis emerges between ascending and descending angular sweeps at all fields, indicating distinct evolution trends of the domain populations during the field sweeping. These results suggest that the sample exhibits strong in-plane resistivity anisotropy with two-fold symmetry, which is consistent with the expectations of $p$-wave magnetism.
We also perform magnetoresistance (MR) and angular magnetoresistance (AMR) measurements at various temperatures. Interestingly, similar behaviors are observed at 7 K [Figs. 2(c) and 2(d)], where the sample resides in the incommensurate spin-density wave (SDW) state, albeit with minor differences in the hysteresis features. Although a finite angular variation of resistivity is expected in this phase due to the stripe-like magnetic structure, the observed strong in-plane anisotropy is still surprising for a system with perfectly collinear AFM order. This may be attributed to the pronounced structural anisotropy of the stripe-like magnetic sublattice, which could enable the emergence of Néel spin currents along the stripe direction [28].
Figure 3: (a) and (b) Angular dependence of resistance at 9 T below and above $T_{N1}$, respectively. (c) and (d) Temperature dependence of the hysteretic angle $\delta\theta$ in AMR at 9 T and maximum hysteretic resistance $\Delta R$ in MR, respectively.
Figures 3(a) and 3(b) display the AMR at 9 T below and above $T_{N1}$, respectively. The squared two-fold angular dependence persists up to 8 K and vanishes near $T_{N1}$. Above $T_{N1}$, the AMR follows a conventional trigonometric form that originates from the Lorentz effect. Although the hysteresis behavior shows no significant difference in different phases, $\Delta\theta$ at 9 T, defined as the angular difference between ascending and descending angular scans, decreases initially with increasing temperature, peaks at $T_{N2}$ and decreases rapidly to zero approaching $T_{N1}$ [Fig. 3(c)]. Intriguingly, while $\Delta R/R_{AVE}$, defined as the ratio of the difference between maximum and minimum resistance at the largest MR hysteresis to their average value, shows a similar temperature trend, its peak lies between $T_{N1}$ and $T_{N2}$ with a minimum at $T_{N2}$ [Fig. 3(d)]. These results suggest complex domain dynamics modulated by magnetic field amplitude and direction.
Figure 4: (a) Temperature dependence of anisotropic ratio $\zeta$(0T) (as defined in the text). Both $\zeta_{FT}$(0T) and $\zeta_{FT}^{ZFW}$(0T) are measured at zero field. However, the former is measured after 9-T field training at each temperature, while the latter is obtained after 9-T field training at 1.8 K followed by zero-field warming. (b) Temperature dependence of $\zeta$(9T). (c) Control of $\zeta$(0T) at 1.8 K with magnetic field alternating along $x$ and $y$ directions.
To quantitatively study the resistivity anisotropy, we define the dimensionless anisotropic ratio $\zeta$ as $(R_{y}-R_{x})/R_{x}$, where the subscripts denote the field orientation parallel to $y$ and $x$ axis, respectively. Due to hysteresis in MR, $\zeta$ clearly depends on field-application history. To obtain $\zeta$ at zero field, we employ the following field-training process: the sample is zero-field cooled to the target temperature, then subjected to 9 T, and finally returned to zero field. Figure 4(a) shows the temperature dependence of $\zeta_{FT}$(0T), which decreases initially with increasing temperature, exhibits a hump slightly above $T_{N2}$, and approaches zero above $T_{N1}$. $\zeta_{FT}^{ZFW}$(0T), which is field-trained at 1.8 K followed by zero-field warming, shows similar values at low temperatures but exhibits no hump around $T_{N2}$ and decreases rapidly to zero when approaching $T_{N1}$ [Fig. 4(a)]. While $T_{N2}$ is discernible in $\zeta$(0T), it becomes nearly indistinguishable in $\zeta$(9T) as shown in Fig. 4(b).
The nonvolatile, field-driven domain selection and resistivity switching in CeNiAsO demonstrate its potential for field-controlled memories. We therefore implement a writing protocol using pulse-like in-plane magnetic fields to deterministically train the domains. The written states are then read out by exploiting the resistance states associated with the domain configuration [Fig. 4(c)]. At 1.8 K, reversible switching between high- and low-resistance states is achieved by applying alternating magnetic field pulses along the x- and y-directions. This write-read cycle can be repeated over multiple iterations without degradation in resistance contrast, demonstrating robustness and non-volatility.
Theoretical studies of CeNiAsO predict that large in-plane resistance anisotropy serves as a direct experimental signature of its $p$-wave magnetism [7]. Our results demonstrate that the anisotropy of the conductivity, $(\sigma_{yy}-\sigma_{xx})/(\sigma_{xx}+\sigma_{yy})$ as defined in Ref. [7], is approximately 0.15 at 1.8 K, which quantitatively agrees with the theoretically calculated value and thus provides strong evidence for $p$-wave magnetism.
Notably, the resistivity anisotropy is barely observable in the zero-field-cooled (ZFC) process due to the tetragonal lattice symmetry, but emerges clearly after magnetic field training, as described above. This behavior is well explained by the random distribution of magnetic domains in the ZFC state and the subsequent field-induced domain selection process. This observation has significant implications for AFM spintronics: It is widely believed that controlling antiferromagnets with an external magnetic field is extremely difficult below the spin-flop transition. However, our work demonstrates that the magnetic field can effectively manipulate the relative populations of magnetic domains, leading to a pronounced and nonvolatile resistance change. The observed anisotropy (>> 30%) implies a large On/Off ratio comparable to that of conventional spintronic devices based on the giant magnetoresistance (GMR) effect — yet achieved without the need for complex multilayer structures.
In conclusion, using high-quality single crystals of the $p$-wave magnet candidate CeNiAsO, we confirm that the ground state with the noncollinear Néel order exhibits a large two-fold in-plane resistivity anisotropy. This anisotropy is consistent with theoretical predictions associated with $p$-wave magnetism. Crucially, this effect is only observed after applying an in-plane training magnetic field to select a preferred domain population, revealing a pronounced resistivity anisotropy with a ratio up to 35% and highlighting its promising potential for spintronic applications. Similar resistivity anisotropy emerges in the fully collinear SDW phase at higher temperatures, implying that field-controlled domain selection is also feasible in conventional collinear antiferromagnets. The nonvolatile nature of this field-driven domain selection, coupled with reversible resistivity switching, establishes a pathway toward field-controlled memory devices. Our work not only provides experimental evidence for the $p$-wave magnetic character of CeNiAsO but also opens a new route toward high-performance AFM spintronics based on the manipulation of magnetic domain populations.
S. L. acknowledges helpful discussions with Prof. Quansheng Wu and Prof. Guoqiang Yu. This work is supported by the National Key Research and Development Program of China (Grants No. 2022YFA1403400, No. 2021YFA1400400, No. 2022YFA1403800), the National Natural Science Foundation of China (Grants Nos. 12274411, 12241405, and 52250418), the CAS Superconducting Research Project (Grant No. SCZX-0101), the Basic Research Program of the Chinese Academy of Sciences Based on Major Scientific Infrastructures (Grant No. JZHKYPT-2021-08), and the CAS Project for Young Scientists in Basic Research (Grant No. YSBR-084). This study was also partially supported by Synergetic Extreme Condition User Facility (SECUF).